The zoo sold tickets at a special rate on Mother’s Day weekend. 8,400 tickets were sold over the course of Saturday and Sunday. Adult tickets were reduced to $6.00 and children’s tickets to $4.00. Ticket sales for the weekend totaled $40,000. How many adult tickets and children’s tickets were sold over the weekend?

Answer by lwsshak3(11628)

The zoo sold tickets at a special rate on Mother’s Day weekend. 8,400 tickets were sold over the course of Saturday and Sunday. Adult tickets were reduced to $6.00 and children’s tickets to $4.00. Ticket sales for the weekend totaled $40,000. How many adult tickets and children’s tickets were sold over the weekend?
***
let x=number of adult tickets sold
8400-x=number of children’s tickets sold
..
$6x+$4(8400-x)=40000
6x+33600-4x=40000
2x=6400
x=3200
8400-x=5200
number of adult tickets sold=3200
number of children’s tickets sold=5200